Using AX=BY+C
solve and graph:
x-2y=4
: Using AX=BY+C
solve and graph:
x-2y=4

Answer by jim_thompson5910(9917)


x-2y=4 Start with the given equation.


-2y=4-x Subtract x from both sides.


-2y=-x+4 Rearrange the terms.


y=(-x+4)/(-2) Divide both sides by -2 to isolate y.


y=((-1)/(-2))x+(4)/(-2) Break up the fraction.


y=(1/2)x-2 Reduce.


Looking at y=(1/2)x-2 we can see that the equation is in slope-intercept form y=mx+b where the slope is m=1/2 and the y-intercept is b=-2


Since b=-2 this tells us that the y-intercept is .Remember the y-intercept is the point where the graph intersects with the y-axis

So we have one point

drawing(500,500,-10,10,-10,10,<BR>
  grid(1),<BR>
  blue(circle(0,-2,.1)),<BR>
  blue(circle(0,-2,.12)),<BR>
  blue(circle(0,-2,.15))<BR>
)


Now since the slope is comprised of the "rise" over the "run" this means
slope=rise/run

Also, because the slope is 1/2, this means:

rise/run=1/2


which shows us that the rise is 1 and the run is 2. This means that to go from point to point, we can go up 1 and over 2



So starting at , go up 1 unit
drawing(500,500,-10,10,-10,10,<BR>
  grid(1),<BR>
  blue(circle(0,-2,.1)),<BR>
  blue(circle(0,-2,.12)),<BR>
  blue(circle(0,-2,.15)),<BR>
  blue(arc(0,-2+(1/2),2,1,90,270))<BR>
)

and to the right 2 units to get to the next point
drawing(500,500,-10,10,-10,10,<BR>
  grid(1),<BR>
  blue(circle(0,-2,.1)),<BR>
  blue(circle(0,-2,.12)),<BR>
  blue(circle(0,-2,.15)),<BR>
  blue(circle(2,-1,.15,1.5)),<BR>
  blue(circle(2,-1,.1,1.5)),<BR>
  blue(arc(0,-2+(1/2),2,1,90,270)),<BR>
  blue(arc((2/2),-1,2,2, 180,360))<BR>
)


Now draw a line through these points to graph y=(1/2)x-2

drawing(500,500,-10,10,-10,10,<BR>
  grid(1),<BR>
  graph(500,500,-10,10,-10,10,(1/2)x-2),<BR>
  blue(circle(0,-2,.1)),<BR>
  blue(circle(0,-2,.12)),<BR>
  blue(circle(0,-2,.15)),<BR>
  blue(circle(2,-1,.15,1.5)),<BR>
  blue(circle(2,-1,.1,1.5)),<BR>
  blue(arc(0,-2+(1/2),2,1,90,270)),<BR>
  blue(arc((2/2),-1,2,2, 180,360))<BR>
) So this is the graph of y=(1/2)x-2 through the points and